Kindly answer this as soon as possible. Urgent Elaborate each step.


Show that Euclidean space and unitary space are not compact. Explain each step. 

1
Expert's answer
2021-10-13T17:04:09-0400

Euclidian space sometimes called the cartesian space is the space of n-tuples of real numbers "(r_1, r_2,...r_n)". It is denoted by "\reals^n"

recall that space x is said to be compact if each of its open cover has a finite subcover.

Euclidian space is not compact:

Reason: Take "0=(0,0,0,...0)\in \reals^n"

consider an open ball of radius R around o for each "k\in \N." denote it by "B_k(0,k)."

"\reals^n \subseteq \displaystyle\cup_{k=1}^{\infin} B_k(0,k)." But it has no finite subcovering corresponding to the covering "\{B_k(0,k)\}K\in\N."

Hence, it is not compact.

Recall that a unitary space is n-dimensional complex linear space on which there is an inner product.

recall that for any "r,y\in t"

"d(r,y)=\sqrt{<r-y,r-y>}"

similar to euclidian space, we may take any point "r\space t\space \in" and construct an open cover at "E\subseteq\displaystyle\cup_{k=1}^{\infin}(n,k)," having no finite subcover.

This gives that unitary space is also not compact.